OpenStudy (anonymous):

In a class 70% passed in mathematics and 80 % passed in English. Together 40% failed in maths and English in a class of 133 members. How many students passed in both the subjects

OpenStudy (acacia):

= 0.7(133) + 0.8(133) - 2(0.4[133]) = 93 + 106 - 2(53) = 199 - 106 = 93

OpenStudy (acacia):

Is that one of your options?

OpenStudy (anonymous):

@acacia could u explain me with venn diagrams becoz the answer which u gave could be found in yahoo answers

OpenStudy (acacia):

Yeah hold on

OpenStudy (texaschic101):

If 70% passed in math, that means 30% failed. If 80% passed in english, that means 20% failed. That totals 50% failed. So how can 40% fail ?

OpenStudy (texaschic101):

@skullpatrol ...can you take a peek at this.....please

OpenStudy (texaschic101):

@Jhannybean ...can you explain this to me please

OpenStudy (texaschic101):

@SithsAndGiggles ...help

OpenStudy (anonymous):

@texaschic101, it's possible that some students are in math, and some in English, but not both.

OpenStudy (texaschic101):

In total, there is 133 students ....right ? In both classes total ?

OpenStudy (texaschic101):

40% failed in a class of 133. (.40)(133) = 53.2 failed ?

OpenStudy (texaschic101):

anybody ?

OpenStudy (anonymous):

@Suchandrarao, do all 133 students take both classes?

OpenStudy (texaschic101):

That is what I was wondering.....is there a total of 133 students in total ?

OpenStudy (amistre64):



OpenStudy (anonymous):

yes @SithsAndGiggles

OpenStudy (amistre64):

.7(133) = 94 .8(133) = 107 107 94 ---- 201 -133 ----- 68/2 = 34 in AnB

OpenStudy (texaschic101):

so 34 people passed in both subjects ?

OpenStudy (amistre64):

well, i dint account for the failures

OpenStudy (amistre64):

.4(133) = 54 students not acowingspaned for on the outsides

OpenStudy (jhannybean):

We divide the total by 2 since there are only 2 subjects, if there were 3, we'd divide our total by 3?

OpenStudy (texaschic101):

107 + 94 + 34 = 235 people in total ?

OpenStudy (amistre64):

133 -54 ---- 79 people to account for passing i believe

OpenStudy (jhannybean):

Why did you delete,@SithsAndGiggles ?

OpenStudy (amistre64):

deciphering this thing is half the battle .... im not sure if ive read it correctly yet or not tho

OpenStudy (anonymous):

@Jhannybean I'm having my doubts.

OpenStudy (anonymous):

@SithsAndGiggles actually there are no options given here

OpenStudy (amistre64):

im thinking 61 is passed AnB

OpenStudy (anonymous):

I think this is what's going on here: \[\begin{align*}n(Class)&=n(Pass)+n(Fail)\\ 133&=\bigg[n(M_p)+n(E_p)-\color{red}{n(M\cap E_p)}\bigg]+\bigg[\color{blue}{n(M_f)}+\color{blue}{n(E_f)}-n(M\cap E_f)\bigg] \end{align*}\] The red is what you're looking for, but I don't think you can determine it without first figuring out the blues. (By the way, \(n(\cdots)\) is the "number of elements in the set \(\cdots\).)

OpenStudy (texaschic101):

you can't do 70% of 133 or 30% of 133 because there is more then 133 students...right ?

OpenStudy (amistre64):



OpenStudy (texaschic101):

We don't know the total number of students...there is more then 133....right ?

OpenStudy (amistre64):

.3k + .2k - .4k = k (.1) failed math or english

OpenStudy (texaschic101):

10% of what failed math or english ?

OpenStudy (amistre64):

"in a class of 133 members" so k = 133

OpenStudy (anonymous):

\[\begin{align*}133&=\bigg[93+106-\color{red}{n(M\cap E_p)}\bigg]+\bigg[\color{blue}{n(M_f)}+\color{blue}{n(E_f)}-53\bigg]\\ 13&=\color{red}{n(M\cap E_p)}-\color{blue}{n(M_f)}-\color{blue}{n(E_f)} \end{align*}\]

OpenStudy (amistre64):

is the question asking: how students passed math and english together? or is it asking how many students passed math? How many students passed english?

OpenStudy (amistre64):

if 10% (14 students): (failed math OR failed english) -(failed math OR failed english) = (passed math AND passed english) = 133 - 14

OpenStudy (zarkon):

if we let M' be the event that we fail math and E' we fail english...is the problem sayin \(P(M'\cap E')=.4\) ?

OpenStudy (texaschic101):

where are you getting the 14 students ?

OpenStudy (anonymous):

I think we need the number of students passed in maths and english i mean if am a student then i should be pass maths as well as english

OpenStudy (amistre64):

from the setup that 30% failed math, and 20% failed english that gives us a sum total of 50% failed both MandE, but it is stated that 40% failed MandE P(fM or fE) = .3k + .2k - .4k = .1k 10% of 133 is 14 kids

OpenStudy (texaschic101):

oh....you rounded up ?

OpenStudy (amistre64):

well, if you rnd down you get less than 10% so you always want to rnd up to cover it

OpenStudy (zarkon):

\[.4=P(M'\cap E')\le P(M')=.3\]

OpenStudy (jhannybean):

In the chart you made, @amistre64 , what are the 1's signifying?

OpenStudy (texaschic101):

total of 100%...I think

OpenStudy (amistre64):

that 100% of the students took math; and 100% of the students took english

OpenStudy (amistre64):

or maybe that 100% of the students the subject are accounted for

OpenStudy (anonymous):

thats what a class consists of 133 students

OpenStudy (jhannybean):

Oh i see.

OpenStudy (texaschic101):

That can't be the total number of students in both classes total....it doesn't equal

OpenStudy (amistre64):

how so

OpenStudy (anonymous):

oh god this question is making hell out of here

OpenStudy (texaschic101):

passing math : (.70)(133) = 93.1....rnd up 92 passing english : (.80)(133) = 106.4 ...rnd up 107 passing math and english : (.60)(133) = 79.8 ...rnd up 80 107 + 80 + 92 = 279.....that is more then 133

OpenStudy (zarkon):

I'm going to type this again... \[.4=P(\text{Fail math}\cap \text{Fail english})\le P(\text{Fail math})=.3\]

OpenStudy (texaschic101):

you see my avatar face.....this is what my real face looks like when I look at that...lol

OpenStudy (amistre64):

i still claim uncertainty :)

OpenStudy (texaschic101):

I am so lost on this.....I am leaving it up to somebody that knows what they are doing...and that is not me...sorry

OpenStudy (zarkon):

does anyone disagree with what i typed above?

OpenStudy (amistre64):

i dont disagree with it :)

OpenStudy (zarkon):

then we have \(.4\le .3\) which is ridiculous. Therefore the problem is ill posed